OpenStudy (anonymous):

What does the line -4x + 3y = -7 look like?

OpenStudy (phi):

there are a few ways to find out. one way is find two points on the line, plot the two points, and draw a line through those points. to find a point, let x=0 in the equation -4*0 + 3y= -7 and solve for y. we get y = -7/3 = -2 ⅓ another point: let y=0 and solve for x -4x +3*0 = -7 -4x = -7 x= -7/(-4) = 7/4 = 1 ¾ plot (0, -2 ⅓) and (1 ¾ , 0)